OpenStudy (djbreezy):

which equation in point-slope form for the given point and slope? Point:(-2,4); Slope:5 A. y+4=5x-10 B. y+4=5(x+2) C. y-4=5(x+2) D. y-4=5(x-2)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (jeansweets):

Y-4=5(X+2)

OpenStudy (skullpatrol):

$$\Huge y- y_0=m(x-x_0)$$

OpenStudy (djbreezy):

is it D?? @skullpatrol

OpenStudy (djbreezy):

I have 3 more @skullpatrol can u help??

OpenStudy (skullpatrol):

post a new question :-)

OpenStudy (djbreezy):

k

OpenStudy (skullpatrol):

The answer is C

OpenStudy (jeansweets):

You just have to get it in the formula plug it in y-y1=m(X-x1) 4 is your y1 -2 is your x1 5 means slope and m The 2 becomes positive because of the double - - signs